OpenStudy (piercetheveil47):

If y − 3 = 3x, which of the following sets represents possible inputs and outputs of the function, represented as ordered pairs? {(0, 3), (1, 6), (2, 9)} {(3, 0), (6, 1), (9, 2)} {(1, 3), (2, 6), (3, 9)} {(3, 1), (6, 2), (9, 3)}

OpenStudy (piercetheveil47):

@dglicks

OpenStudy (anonymous):

It is the first option. Do you need me to explain why?

OpenStudy (anonymous):

I believe its a..

OpenStudy (piercetheveil47):

no, you don't have to if you don't want to

OpenStudy (piercetheveil47):

:)

OpenStudy (anonymous):

All you need to do is Plug the Y axis into y and the x axis into x like this 0y-3=3(3)

OpenStudy (anonymous):

Actually you are going to want to think of your x values as the input and y as the output. When solving for this problem plug the x values in for x and then solve for y. If y corresponds to the output given, then you know that x is valid.

OpenStudy (anonymous):

so for the first example: (0,3) y-3=3x y-3=3(0) y-3=0 y=3